# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ 6.1 CMMI for symmetric indefinite matrices
$\mathbf{W}^{(1,2)}=\underset{\mathbf{O}\in\mathcal{O}_{d_{+},d_{-}}}{\operatorname{argmin}}\|\hat{\mathbf{X}}^{(1)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}\mathbf{O}-\hat{\mathbf{X}}^{(2)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}\|_{F}.$ (6.3)
However, in contrast to the noiseless case, there is no longer an analytical solution to Eq. (6.3). We thus replace Eq. (6.3) with the unconstrained least squares problem
@@
$\mathbf{W}^{(1,2)}=\underset{\mathbf{O}\in\mathbb{R}^{d\times d}}{\operatorname{argmin}}\|\hat{\mathbf{X}}^{(1)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}\mathbf{O}-\hat{\mathbf{X}}^{(2)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}\|_{F},$
whose solution is once again $\mathbf{W}^{(1,2)}=(\hat{\mathbf{X}}^{(1)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle})^{\dagger}\hat{\mathbf{X}}^{(2)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}$. Given $\mathbf{W}^{(1,2)}$, we estimate $\mathbf{P}_{\mathcal{U}_{1},\mathcal{U}_{2}}$ by $\hat{\mathbf{P}}_{\mathcal{U}_{1},\mathcal{U}_{2}}=\hat{\mathbf{X}}^{(1)}\mathbf{W}^{(1,2)}\mathbf{I}_{d_{+},d_{-}}\hat{\mathbf{X}}^{(2)\top}.$ Extending the above idea to a chain of overlapping submatrices is also straightforward; see Section B.1 in the supplementary material for the detailed algorithm and simulation results.
@@
The following result extends Theorem 2 to the indefinite setting. We note that the main difference in this extension is in the upper bound for $\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}$ and this is due to the fact that the least square transformations $\{(\hat{\mathbf{X}}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle})^{\dagger}\hat{\mathbf{X}}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\}_{1\leq\ell\leq L}$ have spectral norms that can be smaller or larger than $1$, and the accumulated error induced by these transformations need not grow linearly with $L$. Finally if $L=1$ then the bounds in Theorem 4 are almost identical to those in Theorem 1, but with a slightly different definition for $\alpha_{i,j}$.
Theorem 4.
@@
$\vartheta_{i_{\ell-1},i_{\ell}}$ $:=\lambda_{\max}(\mathbf{X}^{(i_{\ell})\top}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\mathbf{X}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}),\quad\theta_{i_{\ell-1},i_{\ell}}:=\lambda_{\min}(\mathbf{X}^{(i_{\ell-1})\top}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\mathbf{X}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}),$
$\alpha_{i_{\ell-1},i_{\ell}}:=$ $\frac{n_{i_{\ell-1},i_{\ell}}\gamma_{i_{\ell-1}}\gamma_{i_{\ell}}}{\theta_{{i_{\ell-1}},{i_{\ell}}}(q_{i_{\ell-1}}n_{i_{\ell-1}}\mu_{i_{\ell-1}})^{1/2}(q_{i_{\ell}}n_{i_{\ell}}\mu_{i_{\ell}})^{1/2}}+\frac{n_{{i_{\ell-1}},{i_{\ell}}}^{1/2}}{\theta_{{i_{\ell-1}},{i_{\ell}}}^{1/2}}\Bigl{(}\frac{\gamma_{i_{\ell-1}}^{2}}{q_{i_{\ell-1}}n_{i_{\ell-1}}\mu_{i_{\ell-1}}^{3/2}}+\frac{\vartheta_{{i_{\ell-1}},{i_{\ell}}}^{1/2}\gamma_{i_{\ell}}^{2}}{\theta_{{i_{\ell-1}},{i_{\ell}}}^{1/2}q_{i_{\ell}}n_{i_{\ell}}\mu_{i_{\ell}}^{3/2}}\Bigr{)}$
@@
$\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}=\mathbf{E}^{(i_{0})}\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{0}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{0}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}+\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{L}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{E}^{(i_{L})}+\mathbf{R}^{(i_{0},i_{L})}+\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})},$
@@
6.2 CMMI for asymmetric matrices
Data integration for asymmetric matrices has many applications including genomic data integration (Maneck et al., 2011; Cai et al., 2016), single-cell data integration (Stuart et al., 2019; Ma et al., 2024). Suppose $\mathbf{P}\in\mathbb{R}^{N\times M}$ is a low-rank matrix. Let $d$ be the rank of $\mathbf{P}$, and write the singular decomposition of $\mathbf{P}$ as $\mathbf{P}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top},$ where $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ is a diagonal matrix whose diagonal entries are composed of the singular values of $\mathbf{P}$ in a descending order, and orthonormal columns of $\mathbf{U}\in\mathbb{R}^{N\times d}$ and $\mathbf{V}\in\mathbb{R}^{M\times d}$ constitute the corresponding left and right singular vectors, respectively. The left and right latent position matrices associated to the entities can be represented by $\mathbf{X}=\mathbf{U}\bm{\Sigma}^{1/2}\in\mathbb{R}^{N\times d}$ and $\mathbf{Y}=\mathbf{V}\bm{\Sigma}^{1/2}\in\mathbb{R}^{M\times d}$, respectively. For the $i$th source we denote the index set of the entities for rows and columns by $\mathcal{U}_{i}\subseteq[N]$ and $\mathcal{V}_{i}\subseteq[M]$, and let $\mathbf{P}^{(i)}=\mathbf{P}_{\mathcal{U}_{i},\mathcal{V}_{i}}=\mathbf{U}_{\mathcal{U}_{i}}\bm{\Sigma}\mathbf{V}_{\mathcal{V}_{i}}^{\top}=\mathbf{X}_{\mathcal{U}_{i}}\mathbf{Y}_{\mathcal{V}_{i}}^{\top}.$ For each noisily observed realization $\mathbf{A}^{(i)}$ of $\mathbf{P}^{(i)}$, we obtain the estimated left latent positions $\hat{\mathbf{X}}^{(i)}=\hat{\mathbf{U}}^{(i)}(\hat{\bm{\Sigma}}^{(i)})^{1/2}$ for entities in $\mathcal{U}_{i}$ and right latent positions $\hat{\mathbf{Y}}^{(i)}=\hat{\mathbf{V}}^{(i)}(\hat{\bm{\Sigma}}^{(i)})^{1/2}$ for entities in $\mathcal{V}_{i}$.
Figure 10: A pair of overlapping observed submatrices of an asymmetric/rectangular matrix
Let $\mathbf{P}^{(1)}$ and $\mathbf{P}^{(2)}$ be two overlapping submatrices shown in Figure 10 without noise or missing entries. Suppose $\mathrm{rank}(\mathbf{P}^{(1)})=\mathrm{rank}(\mathbf{P}^{(2)})=\mathrm{rank}(\mathbf{X}_{\mathcal{U}_{1}\cap\mathcal{U}_{2}})=\mathrm{rank}(\mathbf{Y}_{\mathcal{V}_{1}\cap\mathcal{V}_{2}})=d$. Now

### v4

@@ Appendix B Integration of Multiple Matrices with Complex Connectivity
    $\mathbf{W}^{(i,j)}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{O}-\widehat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|_{F}.$
  3. 3.
    Choose one of the vertex denoted by $i_{\star}$} (for example, $i_{\star}=1$), and let $\widetilde{\mathbf{X}}^{(i_{\star})}=\widehat{\mathbf{X}}^{(i_{\star})}$.
  4. 4.
[…]
Step 3 Obtaining the holistic latent position estimate $\widetilde{\mathbf{X}}\in\mathbb{R}^{N\times d}$: For each $\ell\in[N]$, compute $S=\sum_{\ell\in\mathcal{U}_{i}}c_{i}^{-2}$, and compute the holistic estimated latent position as
$\widetilde{\mathbf{X}}_{\ell}=\sum_{\ell\in\mathcal{U}_{i}}(c_{i}^{-2}/S)\widehat{\mathbf{X}}^{(i)}_{\langle\ell\rangle}.$
$\widehat{\mathbf{P}}=\widetilde{\mathbf{X}}\widetilde{\mathbf{X}}^{\top}$.
Our theoretical results and the analysis of Algorithm 1 can be naturally extended to Algorithm B.1, albeit at the cost of more involved expressions and notations. In particular, as the holistic latent position estimate for some entities is computed as a weighted average of individual latent position estimates, the (entrywise) estimation error of the final $\widehat{\mathbf{P}}$ is a weighted average of the errors from the individual estimates. More specifically, recall that the dominant term in the error for each individual estimate is $\mathbf{E}^{({i_{0}})}\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{0}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{0}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}+\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{L}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{E}^{({i_{L}})}$. The dominant term in the error for the holistic latent position estimate is then a weighted average of multiple independent noise terms corresponding to different $(\mathbf{E}^{(i_{0})},\mathbf{E}^{(i_{L})})$ pairs. As a result, the holistic estimate, which aggregates information from multiple blocks, can yield significantly reduced error compared to individual estimates and lead to more accurate recovery. Furthermore, as the dominant error term is a weighted sum of independent terms, one can derive an entrywise normal approximation for the estimation error of $\widehat{\mathbf{P}}$ using a similar analysis to that in Theorem 3 (but with a more complicated expression for the variance $\widetilde{\sigma}_{s,t}^{2}$)
Compared with BONMI in Zhou et al. (2023), which integrates overlapping submatrices pairwise and selects the estimate for each unobserved entity based on the pair with the lowest sum of estimated noise levels, the generalized CMMI algorithm in Algorithm B.1 provides a more refined strategy for multiple matrix integration. First, CMMI aligns all connected submatrices (not just directly overlapping ones), which can significantly expand the set of recoverable entries. Moreover, CMMI simultaneously leverages all connected submatrices by jointly aligning them and computing a weighted average of the estimated latent positions. In contrast, BONMI utilizes only a single selected pair of submatrices for each entity and therefore does not fully exploit the information available from all observed submatrices.
Compared with the sequential integration approach SPSMC in Bishop and Yu (2014), CMMI also provides a more effective integration strategy. CMMI aligns the estimated latent positions from all submatrices more optimally by jointly considering all overlapping pairs, whereas SPSMC is constrained by its sequential structure, and the problem of determining an effective integration order is not addressed in Bishop and Yu (2014). Furthermore, CMMI aggregates the aligned embeddings by incorporating information from all available submatrices, leading to more accurate estimates. In contrast, SPSMC assigns each entity a latent position based solely on the first submatrix in which it appears in the sequence, and thus also does not to fully leverage all available information.
B.1 Real data experiment: MEDLINE co-occurrences
We note that the analysis of the MEDLINE co-occurrence dataset in Section 5 is based on a synthetic scenario where some of the observed entries are held out. While this might lead to a somewhat artificial use case, it is nevertheless intentional as we can then evaluate the performance of CMMI for simple chains (as implemented in Algorithm 1). We now consider a more realistic data integration problem for the MEDLINE data. More specifically, if we partition the citations by year then each year tends to feature a different set of frequently occurring clinical concepts. The PMIs computed between the high-frequency concepts in a given year are likely to be less noisy compared to those involving rarely occurring concepts. We now demonstrate that, by extracting the PMIs for top-frequency concepts in each year and then integrating them using CMMI (or other matrix integration algorithms), we recover more accurate co-occurrence relationships between the clinical concepts than the PMIs computed directly from data aggregated across all the years.
We consider MEDLINE co-occurrence data from the years 1993 to 2022. For each year, we extract a PMI submatrix based on the co-occurrence of the top 1000 most frequent clinical concepts. Given a number of observed years $K$ (it is also the number of observed PMI submatrices), we aim to integrate these submatrices. In this experiment, we always select the most temporally distant years for the PMI submatrix integration task. For example, for $K=2$, we integrate the PMI submatrices corresponding to the years 1993 and 2022. These two $1000\times 1000$ PMI submatrices together involve a total of $N=1540$ unique clinical concepts. Our goal is to recover the unobserved entries (approximately 25%) in the resulting $1540\times 1540$ PMI matrix.
When applying CMMI in Algorithm B.1, we determine the embedding dimension $d$ by first applying the automatic dimensionality selection procedure from Zhu and Ghodsi (2006) to each observed submatrix, and then selecting the largest resulting value as $d$ to retain sufficient information. For example, when $K=2$, the procedure yields dimensions $12$ and $16$ for the two submatrices, and we set $d=16$ for CMMI. For FALS, we use the same dimension $d$ as in CMMI, while for $GSR$, we always fix $d=3$ to avoid excessive computational cost. In addition, for low-rank matrix completion algorithms other than CMMI, we first construct a global matrix by merging all observed submatrices, where entries appearing in multiple submatrices are averaged, and then apply the matrix completion algorithm to this aggregated matrix.
We refer to the pre-trained clinical concept embeddings from Beam et al. (2020), learned from massive sources of multimodal medical data, to evaluate how well the algorithms recover the unobserved entries. Given pre-trained embedding vectors $\bm{v}_{1},\dots,\bm{v}_{N}$ for $N$ clinical concepts, we construct a similarity matrix $\mathbf{P}\in\mathbb{R}^{N\times N}$ where each entry $\mathbf{P}_{ij}$ is the cosine similarity between $\bm{v}_{i}$ and $\bm{v}_{j}$. We then measure the similarities between the estimated PMIs for the unobserved entries with corresponding entries in $\mathbf{P}$ in terms of the Spearman’s rank correlation.
We vary $K$ from $2$ to $15$ to compare the performance of different algorithms, and the results are summarized in Figure B.2. Note that the baseline, shown as a dashed black line in the left panel of Figure B.2, represents the performance obtained by directly computing PMIs from the co-occurrence data aggregated across the selected years. Figure B.2 shows that integrating per-year PMIs can yield more faithful co-occurrence relationships than directly computing PMIs from aggregated data, especially when only a few years are observed. For example, when $K=2$, the baseline achieves only $0.048$, while CMMI reaches $0.273$. And Figure B.2 also shows that, compared to other matrix integration algorithms, CMMI has the highest accuracy while maintaining significant advantages in computational efficiency.
Figure B.2: The left panel reports empirical estimates of Spearman’s rank correlations for CMMI and other low-rank matrix completion algorithms as we vary $K$ from $2$ to $15$. Note that we always use the most temporally distant $K$ years for the PMI submatrix integration task. For example, when $K=2$, we use the years 1993 and 2022; when $K=3$, we use the years 2022, 2008, and 1994; and when $K=15$, we use the years 2022, 2020, …, 1994. On the x-axis, each value of $K$ is followed (in parentheses) by the proportion of unobserved entries to be recovered in the entire matrix. The running time (in log scale) for algorithms is shown in the right panel.
B.2 Computational running time for Sections 4.2, 5, and A.4
In the simulation setting in Section 4.2 and the synthetic settings in Sections 5 and A.4, only a single missing submatrix was of interest, and the basic CMMI was used to impute this submatrix. In contrast, other existing low-rank matrix completion algorithms are designed to impute all missing entries across the entire matrix, which makes direct comparisons of computational time somewhat unfair. With the generalized CMMI, we are now able to impute all unobserved entries, thereby allowing for a fairer comparison of computational performance.
The results presented below are obtained following the same setup as in Sections 4.2,5, and A.4. Instead of using the basic CMMI in Algorithm 1 to impute just a single missing block, we apply the generalized CMMI in Algorithm B.1 to recover all unobserved entries in the entire matrix, and in evaluating algorithm performance, we consider all unobserved entries rather than focusing on just a single block of interest.
Additional experimental results for Section 4.2:
Figure B.3: The left panel reports empirical errors $\|\widehat{\mathbf{P}}_{\text{unobserved}}-\mathbf{P}_{\text{unobserved}}\|_{F}/\|\mathbf{P}_{\text{unobserved}}\|_{F}$, where $\mathbf{P}_{\text{unobserved}}$ and $\widehat{\mathbf{P}}_{\text{unobserved}}$ denote the submatrices of unobserved entries in $\mathbf{P}$ and $\widehat{\mathbf{P}}$, respectively, for CMMI and other matrix completion algorithms as we vary $L\in\{1,2,3,4,7,9,14,19\}$ while fixing $N\approx 2200$, $\breve{p}=0.1$, $q=0.8$, $\sigma=0.5$. The results are averaged over $100$ independent Monte Carlo replicates. Note that the averaged relative $F$-norm errors of IRRSV are $\{0.8,4.4,11.5,15.2,32.7,37.5,202.7,2299.0\}$ and some of these values are too large to be displayed in this panel. The average running time (in log scale) over 100 replicates for algorithms, using 20-core parallel computing and 256 GB memory, is shown in the right panel.
Additional experimental results for Section 5:
Figure B.4: The left panel reports empirical estimates of Spearman’s rank correlations for CMMI and other low-rank matrix completion algorithms as $L$ changes. In particular, we vary $L=\{1,2,3,5\}$ while fixing $n=1000$ and $\breve{p}=0.1.$ The results are averaged over $100$ independent Monte Carlo replicates. The average running time (in log scale) over 100 replicates for algorithms, using 25-core parallel computing and 256 GB memory, is shown in the right panel.
Additional experimental results for Section A.4:
Figure B.5: The left panel reports ARIs for clustering $\widehat{\mathbf{X}}\in\mathbb{R}^{N\times d}$, which consists of the top $d=36$ scaled eigenvectors of the recovered full matrix $\widehat{\mathbf{P}}$, obtained using CMMI and other low-rank matrix completion algorithms. In particular, we vary $L\in\{1,2,3,4\}$ while fixing $n=1000$, $\breve{p}=0.1$, $q=0.8$. The results are averaged over $100$ independent Monte Carlo replicates. The average running time (in log scale) over 100 replicates for algorithms, using 25-core parallel computing and 256 GB memory, is shown in the right panel.
Figures B.3, B.4, and B.5 show that CMMI consistently outperforms other low-rank matrix completion algorithms in recovering the entire matrix, and its computational time remains highly competitive.
Appendix C Algorithms and Simulation Results for Section 6
C.1 Symmetric indefinite matrices integration
Algorithm C.1 CMMI for overlapping submatrices of a symmetric indefinite matrix
  1. 1.
    For each $0\leq\ell\leq L$, obtain the estimated latent positions $\widehat{\mathbf{X}}^{(i_{\ell})}=\widehat{\mathbf{U}}^{(i_{\ell})}|\widehat{\bm{\Lambda}}^{(i_{\ell})}|^{1/2}$.
  2. 2.
@@ C.1 Symmetric indefinite matrices integration
    $\mathbf{W}^{(i_{\ell-1},i_{\ell})}$ $=\underset{\mathbf{O}\in\mathbb{R}^{d\times d}}{\operatorname{argmin}}\|\widehat{\mathbf{X}}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\mathbf{O}-\widehat{\mathbf{X}}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\|_{F}.$
  3. 3.
    Compute $\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}=\widehat{\mathbf{X}}^{(i_{0})}\mathbf{W}^{(i_{0},i_{1})}\mathbf{W}^{(i_{1},i_{2})}\cdots\mathbf{W}^{(i_{L}-1,i_{L})}\mathbf{I}_{d_{+},d_{-}}\widehat{\mathbf{X}}^{(i_{L})\top}$.
[…]
@@
Algorithm C.1 presents a procedure for integration of symmetric but possibly indefinite matrices. We now state an extension of Theorem 2 to this setting. The main difference in this extension is the upper bound for $\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})}$ and this is due to the fact that the least square transformations $(\widehat{\mathbf{X}}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle})^{\dagger}\widehat{\mathbf{X}}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}$ for $1\leq\ell\leq L$ have spectral norms that can be smaller or larger than $1$, and the accumulated error induced by these transformations need not grow linearly with $L$. If $L=1$ then the bounds in Theorem C.1 are almost identical to those in Theorem 1, but with a slightly different definition for $\alpha_{i,j}$.
Theorem C.1.
@@
$\vartheta_{i_{\ell-1},i_{\ell}}$ $=\lambda_{\max}(\mathbf{X}^{(i_{\ell})\top}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\mathbf{X}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}),\quad\theta_{i_{\ell-1},i_{\ell}}=\lambda_{\min}(\mathbf{X}^{(i_{\ell-1})\top}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\mathbf{X}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle})$,
$\alpha_{i_{\ell-1},i_{\ell}}$= $\frac{n_{i_{\ell-1},i_{\ell}}\gamma_{i_{\ell-1}}\gamma_{i_{\ell}}}{\theta_{{i_{\ell-1}},{i_{\ell}}}(q_{i_{\ell-1}}n_{i_{\ell-1}}\mu_{i_{\ell-1}})^{1/2}(q_{i_{\ell}}n_{i_{\ell}}\mu_{i_{\ell}})^{1/2}}+\frac{n_{{i_{\ell-1}},{i_{\ell}}}^{1/2}}{\theta_{{i_{\ell-1}},{i_{\ell}}}^{1/2}}\Bigl{(}\frac{\gamma_{i_{\ell-1}}^{2}}{q_{i_{\ell-1}}n_{i_{\ell-1}}\mu_{i_{\ell-1}}^{3/2}}+\frac{\vartheta_{{i_{\ell-1}},{i_{\ell}}}^{1/2}\gamma_{i_{\ell}}^{2}}{\theta_{{i_{\ell-1}},{i_{\ell}}}^{1/2}q_{i_{\ell}}n_{i_{\ell}}\mu_{i_{\ell}}^{3/2}}\Bigr{)}$
@@
$\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}=\mathbf{E}^{(i_{0})}\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{0}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{0}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}+\mathbf{X}_{\mathcal{U}_{i_{0}}}(\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{X}_{\mathcal{U}_{i_{L}}})^{-1}\mathbf{X}_{\mathcal{U}_{i_{L}}}^{\top}\mathbf{E}^{(i_{L})}+\mathbf{R}^{(i_{0},i_{L})}+\mathbf{S}^{(i_{0},i_{1},\dots,i_{L})},$
@@
We compare the performance of Algorithm C.1 with other matrix completion methods. Consider the setting of Section 4.2, but with $\bm{\Lambda}=\mathrm{diag}(N,\tfrac{1}{2}N,-\tfrac{1}{2}N,-N)$, and thus $d_{+}=d_{-}=2$. Figure C.1 shows the relative $F$-norm estimation error results for CMMI against other matrix completion algorithms.
Figure C.1: The left panel reports relative $F$-norm errors $\|\widehat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{F}/\|\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{F}$ for CMMI and other low-rank matrix completion algorithms as $L$ changes. In particular, we vary $L=\{1,2,3,4,7,9,14,19\}$ while fixing $N\approx 2200$, $\breve{p}=0.1$, $q=0.8$, $\sigma=0.5$. The results are averaged over $100$ independent Monte Carlo replicates. Note that the averaged relative $F$-norm errors of IRRSV are $\{0.4,1.9,2.6,3.8,11.0,70.4,469.6,305.0\}$ and some of these values are too large to be displayed in this panel. The average running time (in log scale) over 100 replicates for algorithms, using 20-core parallel computing and 256 GB memory, is shown in the right panel.
C.2 Asymmetric matrices integration
We provide detailed derivations of our idea for asymmetric matrices here. We first consider the noiseless case to illustrate the idea. Let $\mathbf{P}^{(1)}$ and $\mathbf{P}^{(2)}$ be two overlapping submatrices shown in the left panel of Figure 7 without noise or missing entries. Suppose $\mathrm{rk}(\mathbf{P}^{(1)})=\mathrm{rk}(\mathbf{P}^{(2)})=\mathrm{rk}(\mathbf{X}_{\mathcal{U}_{1}\cap\mathcal{U}_{2}})=\mathrm{rk}(\mathbf{Y}_{\mathcal{V}_{1}\cap\mathcal{V}_{2}})=d$. Now